OpenStudy (steparti1):

Li-Mei thinks she can save money on electricity by changing all her bulbs from incandescent bulbs to fluorescent bulbs. Based on the following energy cost table, how much can she save each year if she replaces 8 bulbs? Incandescent bulb - 150 hours monthly use $.05 per hour Fluorescent bulb - 150 hours monthly use $.02 per hour a. $36 b. $432 c. $54 d. $1,008

OpenStudy (anonymous):

Both light bulbs are used 150 hours a month. There are 12 months in a year, so both lightbulbs are used 12*150=1800 hours per year. The difference in cost per hour is 0.05-0.02=0.03. Switching lightbulbs saves $0.03 an hour. Thus, 1800 hours*$0.03 cents saved per hour=$54 saved per year. This is for one lightbulb. The problem says that Li-Mei replaced 8 bulbs. 8*54=$432 total saved.